OpenStudy (rulerlol113):

Express 3a 2b -1 with positive exponents. this is annoying i dont understand please help me

OpenStudy (sleepyjess):

Hello ^_^ Welcome to OpenStudy! Could you type the equation using the Equation Editor button?

OpenStudy (rulerlol113):

\[Express 3a ^{2} b ^{-1} with positive exponents.\]

pooja195 (pooja195):

\[\huge~\rm~\bf~3a^2b^{-1}\]

OpenStudy (rulerlol113):

i need to know how to make that positive it shows a fraction as the answer or maybe its a divison thing

pooja195 (pooja195):

@sleepyjess You may continue helping, I wrote out the problem with the equation tool

OpenStudy (sleepyjess):

Anything with a negative exponent is 1/term^exponent. So if we had \(b^{-8}\), that would turn in to \(\dfrac{1}{b^8}\)

OpenStudy (rulerlol113):

thank you sleepyjess

OpenStudy (sleepyjess):

Here, we have \(b^{-1}\), so what do you think that becomes?

OpenStudy (rulerlol113):

\[\frac{ 1 }{ b ^{1} }\]

OpenStudy (sleepyjess):

Yep :)

OpenStudy (rulerlol113):

what about the rest?

OpenStudy (sleepyjess):

And then we just plug that back into the equation, which becomes \(3a^2(\dfrac 1b)\), which can actually be further simplified to \(\dfrac{3a^2}{b}\)

OpenStudy (rulerlol113):

thank you so much

OpenStudy (sleepyjess):

No problem :)